# Sensor-fusion based Prognostics Framework for Complex Engineering Systems Exhibiting Multiple Failure Modes

### v2

@@ 4 Offline Sensor Selection Algorithm
We consider a scenario where a system is monitored by $P$ sensors. We assume there exists a historical dataset comprised of $P$-dimensional degradation-based sensor signals from monitoring $N$ similar systems. We denote $s_{i,p}(t)$ as the signal data from sensor $p\in\{1,...,P\}$ for system $i\in\{1,...,N\}$ observed at time $t\in[0,T]$, where $T=min(\{TTF_{i}\}_{i=1}^{N})$. By truncating all signals to the minimum failure time, we ensure that all signals in the training set have the same time domain and that they are all included when modeling. For this framework, we assume that the total number of failure modes, $K$ is known, but that the actual failure modes in the training dataset are unknown. This poses a challenge for CA-FPCA, which requires observed covariates for computation. To determine these covariates, we perform a sensor-wise clustering of the training samples. For each sensor, we perform FPCA for feature extraction and then use K-means clustering with $K$ clusters to label the signals from that sensor. Once the sensor-wise clustering is complete, the next step involves feature extraction by applying CA-FPCA and then regressing these features against the time-to-failure to identify the most informative subset of sensors for each failure mode. The next two subsections discuss details about CA-FPCA and the penalized Mixture of Gaussian regression model (MGR-ASGL).
@@ 4.1 Feature Extraction: CA-FPCA
where $v_{i,p}^{k}(t)$ is a smooth random function, and $\epsilon_{i,p}(t)$ are assumed to be independently and identically distributed (i.i.d.) errors with mean zero and variance $\varsigma_{p}^{2}$. $v_{i,p}^{k}(t)$ and $\epsilon_{i,p}(t)$ are assumed to be independent of each other. The mean and covariance functions of $v_{i,p}^{k}(t)$ are given by $\mu_{p}^{k}(t)$ and $\mathbb{C}_{p}^{k}(t,t^{\prime})$, respectively. Using Mercer’s theorem [35], the covariance can be decomposed as follows, $\mathbb{C}_{p}^{k}(t,t^{\prime})=\sum_{m=1}^{\infty}\lambda_{m,p}^{k}\phi_{m,p}^{k}(t)\phi_{m,p}^{k}(t^{\prime})$, where $\{\phi_{m,p}^{k}(t)\}_{m=1}^{\infty}$ are the orthogonal eigenfunctions of $\mathbb{C}_{p}^{k}(t,t^{\prime})$ and $\lambda_{1,p}^{k}\geq...\geq\lambda_{m,p}^{k}\geq...$ are the corresponding eigenvalues. By projecting the mean-subtracted signal data onto the eigenfunctions, we can represent the signals by equation (2).
@@
To ensure all features are scaled equally, we use the standardized versions of the CA-FPC-scores: $x_{i,m}^{p}=\frac{\sum_{k=1}^{K}\xi_{i,m,p}^{k}\mathit{\mathbb{I}}(i\in\mathcal{I}_{p}^{k})}{\sqrt{\frac{1}{N-1}\sum_{k=1}^{r_{p}^{*}}N_{p}^{k}\lambda_{m,p}^{k}}}$, where $N_{p}^{k}$ is the number of signals assigned to cluster $k$ for sensor $p$. We extract $\boldsymbol{x}_{i,p}=(x_{i,1}^{p},...,x_{i,q_{p}}^{p})^{T}$ for sensor $p$ of system $i$,. We then utilize this $\ boldsymbol{x}_{i,p}$’s as predictors for fitting the MGR-ASGL model.
@@ 4.2 Optimization of the MGR-ASGL model
To perform sensor selection, we seek to model the distribution of $\ln{TTF}$ conditioned on the observed set of features. In what follows, we note that values of $\ln{TTF}$ are standardized. Our modeling approach assumes that the failure times are conditionally independent given the observed features. Let $y_{i}$ denote the random variable corresponding to the $\ln{TTF}$ of the system $i$. We consider $\boldsymbol{Z}_{i}$ to denote a “one-hot” random vector of length $K$, where $\boldsymbol{Z}_{i}[k]=1$ if system $i$ failed due to failure mode $k$, and zero otherwise. For mathematical brevity, we let $\boldsymbol{Y}=\left(y_{1},\ldots,y_{N}\right)$ denote the observed $\ln{TTF}$ and $\mathbb{Z}=\left(\boldsymbol{Z}_{1},\boldsymbol{Z}_{2},\ldots,\boldsymbol{Z}_{N}\right)$ denote the set of random vectors representing the failure modes for each system. In the following subsections, we first assign probability distributions to these random variables to formalize the MGR model. Second, we construct the log-likelihood function for the MGR model. Third, we provide theoretical intuition on the use of the EM algorithm for model fitting. Fourth, we describe the ASGL penalty. Finally, we provide details for implementing the EM algorithm.
4.2.1 Mixture of Gaussian Regression (MGR) model
We assume that the distribution of $y_{i}$ conditioned on the observed features $\boldsymbol{x}_{i,1},\ldots,\boldsymbol{x}_{i,P}$} and $\boldsymbol{Z}_{i}[k]=1$ is Gaussian. If we let $\pi_{1},\ldots,\pi_{K}$, where $\sum_{k=1}^{K}\pi_{k}=1$, denote the mixing coefficients, which represent the probability that $\boldsymbol{Z}_{i}[k]=1$ ($\forall k=\{1,\ldots,K\}$), then equation (3) gives the marginal distribution of $y_{i}$ given only $\{\boldsymbol{x}_{i,1},\ldots,\boldsymbol{x}_{i,p}\}$. This marginal distribution is the MGR model.
$f_{\vartheta}(y_{i}\mid\boldsymbol{x}_{i,1},\ldots,\boldsymbol{x}_{i,P})=\sum_{k=1}^{K}\pi_{k}\frac{1}{\sqrt{2\pi}\sigma_{k}}\times\ \exp\left(-\frac{1}{2\sigma_{k}^{2}}\left(y_{i}-\beta_{0,k}-\sum_{p=1}^{P}\boldsymbol{x}_{i,p}^{T}\boldsymbol{\beta}_{p,k}\right)^{2}\right)$ (3)
4.2.2 Log-Likelihood
[…]
Definition 1.
Definition 2.
@@ 4.2.2 Log-Likelihood
$\ell_{C}(\boldsymbol{\Theta};\boldsymbol{Y},\mathbb{Z})=\sum_{i=1}^{N}\sum_{k=1}^{K}\boldsymbol{Z}_{i}[k]\ln\pi_{k}\left(\frac{\rho_{k}}{\sqrt{2\pi}}\right)\exp\left(-\frac{1}{2}\left(\rho_{k}y_{i}-\psi_{0,k}-\sum_{p=1}^{P}\boldsymbol{x}_{i,p}^{T}\boldsymbol{\psi}_{p,k}\right)^{2}\right)$ (5)
Input : Number of systems $N$, Number of sensors $P$, Unlabeled data from $P$ sensors of all $N$ systems, Maximum iterations $iter_{max}$ for EM algorithm, threshold $\delta$ for regression coefficient’s $l_{2}$-norm
Output : Vector $\boldsymbol{Z}$ indicating failure labels, Set of informative sensors $\mathcal{P}_{k}$ for each failure mode $k$
/* Feature Extraction from all $P$ sensors (CA-FPCA) */
1 for $p=1$ to $P$ do
2       $\mathcal{I}_{k}^{p}\leftarrow$ K-means clustering on FPC-scores;
3       for $k=1$ to $K$ do
4             $\xi_{i,m,p}(k)\leftarrow$ Use $\mathcal{I}_{k}^{p}$ to obtain CA-FPC-scores (Eq. 2);
5
6       end for
7      $x_{i,p}\leftarrow$ Obtain standardized CA-FPC-scores;
8       $q_{p}\leftarrow$ Number of CA-FPC-scores retained for sensor $p$;
9
10 end for
/* MGR-ASGL based EM algorithm */
[…]
12 while $iter\leq iter_{max}$ do
       $\gamma_{i}^{iter}\leftarrow$ Use $x_{i,p}$ in Eq. (7) for all $i$; ;
        // E-step
       $\Pi^{iter}\leftarrow$ Update using Eq. (9); ;
        // M-step for $\Pi$
13       $\boldsymbol{\rho}^{iter},\Psi^{iter}\leftarrow$ Solve Eq. (8) using $\sum_{p=1}^{P}\sqrt{q_{p}}$ and $\gamma_{i}^{iter}$;
14       $iter\leftarrow iter+1$;
15
16 end while
[…]
18 $\boldsymbol{Z}_{i}[\operatorname*{arg\,max}_{k}{\gamma_{i}^{iter_{max}}}]\leftarrow 1$ for all $i$;
19 $\boldsymbol{Z}_{i}[\sim\operatorname*{arg\,max}_{k}{\gamma_{i}^{iter_{max}}}]\leftarrow 0$ for all $i$;
20 foreach $p$ and $k$ do
21       if $\lVert\psi^{iter_{max}}_{p,k}\rVert_{2}\geq\delta$ then
             Append $p$ to $\mathcal{P}_{k}$ ;
              // Significant sensors for failure mode $k$
22            ;
23
24       end if
25
26 end foreach
Algorithm 1 Offline Sensor Selection Algorithm
Lemma 1.
Proof.
@@ 18 $\boldsymbol{Z}_{i}[\operatorname*{arg\,max}_{k}{\gamma_{i}^{iter_{max}}}]\leftarrow 1$ for all $i$;
Since $\ln(.)$ is concave, from Jensen’s inequality we can write that,
and thus,
For every iteration of the EM algorithm, the E-step calculates the conditional expectation of the CDLL given the observed failure times, which we denote as $\mathbb{E}_{\boldsymbol{\Theta}_{\text{old}}}[\ell_{C}(\boldsymbol{\Theta};\boldsymbol{Y},\mathbb{Z})|\boldsymbol{Y}]$. In the M-step, we find the parameters that maximize the conditional expectation (see equation (8)). The parameter estimation equations of the EM algorithm are provided in III.B.5.
4.2.3 Adaptive Sparse Group Lasso (ASGL) penalty
The inclusion of features from non-informative sensors can negatively impact the accuracy of the RUL prediction. To address this issue, we penalize the negative-scaled conditional expectation of the CDLL function using the Adaptive Sparse Group Lasso (ASGL) penalty to obtain the following minimization problem given by 6.
$\min_{\boldsymbol{\Theta}}-\frac{1}{N}\mathbb{E}_{\boldsymbol{\Theta}_{\text{old}}}[\ell_{C}(\boldsymbol{\Theta};\boldsymbol{Y},\mathbb{Z})|\boldsymbol{Y}]+\lambda\sum_{k=1}^{K}\pi_{k}\sum_{p=1}^{P}\sqrt{q_{p}}\times\left[(1-\alpha)\left\|\boldsymbol{\psi}_{p,k}\right\|_{2}+\alpha\left\|\boldsymbol{\psi}_{p,k}\right\|_{1}\right]$ (6)
ASGL is an extension of lasso [39] that ensures the coefficients for all features of a non-informative group (sensor) are reduced to zero. In the context of regularization, adaptive refers to the penalty accounting for the number of CA-FPC-scores retained for each sensor i.e., $q_{p}$, and the mixing coefficient for each failure mode i.e., $\pi_{k}$. While the inclusion of $q_{p}$ ensures a sensor is not retained simply by having more coefficients, $\pi_{k}$ accounts for the class imbalance. $\lambda>0$ and $\alpha\in[0,1]$ are the hyperparameters for this model.
4.2.4 Expectation Maximization algorithm
We now describe the EM algorithm used to find a local optimum to the IDLL described. Given an initial estimate of the model parameters $\boldsymbol{\Theta}_{old}$, we perform the E-step and M-step as follows:
E-step: The E-step consists of computing the expectation of CDLL in (6) conditioned on $\boldsymbol{Y}$. Let $\mathcal{Q}\left(\boldsymbol{\Theta}\mid\boldsymbol{\Theta}_{\text{old}}\right)=-\mathbb{E}_{\boldsymbol{\Theta}_{\text{old}}}\left[\ell_{C}(\boldsymbol{\Theta};\boldsymbol{Y},\mathbb{Z})\mid\boldsymbol{Y}\right]$ denote the negative expectation of the CDLL. In order to compute $\mathcal{Q}\left(\boldsymbol{\Theta}\mid\boldsymbol{\Theta}_{\text{old}}\right)$, we need to compute $\gamma_{i,k}=$ $\mathbb{E}_{\boldsymbol{\Theta}_{\text{old }}}\left[Z_{i}[k]\mid\boldsymbol{Y}\right]$, which represents the responsibility of failure mode $k$ to system $i$ and is given by equation (7). For any system $i$, we denote the vector of the responsibilites $\gamma_{i,k}$ across all failure modes as $\boldsymbol{\gamma}_{i}$ such that, $\boldsymbol{\gamma}_{i}=\{\gamma_{i,1},...,\gamma_{i,K}\}$.
$\footnotesize\gamma_{i,k}=\frac{\pi_{k}^{\text{old }}\rho_{k}^{\text{old }}\times\exp\left(-\frac{1}{2}\left(\rho_{k}^{\text{old }}y_{i}-\psi_{0,k}^{\text{old }}-\sum_{p=1}^{P}\boldsymbol{x}_{i,p}^{T}\boldsymbol{\psi}_{p,k}^{\text{old }}\right)^{2}\right)}{\sum_{l=1}^{K}\pi_{l}^{\text{old }}\rho_{l}^{\text{old }}\exp\left(-\frac{1}{2}\left(\rho_{l}^{\text{old }}y_{i}-\psi_{0,l}^{\text{old }}-\sum_{p=1}^{P}\boldsymbol{x}_{i,p}^{T}\boldsymbol{\psi}_{p,l}^{\text{old }}\right)^{2}\right)}\vspace*{-0.5em}$ (7)
M-Step: During the M-step, we update the parameters by solving the optimization problem in (8) with respect to $\boldsymbol{\Theta}$.
$\min_{\boldsymbol{\Theta}}\frac{1}{N}Q(\boldsymbol{\Theta}\mid\boldsymbol{\Theta}_{\text{old}})+\lambda\sum_{k=1}^{K}\pi_{k}\sum_{p=1}^{P}\sqrt{q_{p}}\times\left[(1-\alpha)\left\|\boldsymbol{\psi}_{p,k}\right\|_{2}+\alpha\left\|\boldsymbol{\psi}_{p,k}\right\|_{1}\right]$ (8)
Improvement with respect to $\boldsymbol{\Pi}$: To update the mixing coefficients, we set $\rho_{k}=\rho_{k}^{\text{old}},\boldsymbol{\psi}_{0,k}=\boldsymbol{\psi}_{0,k}^{\text{old }},\boldsymbol{\psi}_{p,k}=\boldsymbol{\psi}_{p,k}^{\text{old }}$. and solve optimization problem (8). We utilize the update strategy from [37]. We start by identifying a feasible solution $\overline{\boldsymbol{\Pi}}^{\text{new}}=\left(\frac{1}{N}\sum_{i=1}^{N}\gamma_{i,1},\ldots,\frac{1}{N}\sum_{i=1}^{N}\gamma_{i,K}\right)^{T}$ and then update the mixing coefficients as described by equation (9), where $t^{\text{new }}\in(0,1]$ is chosen to be the largest value in $\left\{0.1^{k};k=0,1,2,\ldots\right\}$ such that the objective function does not increase.
$\boldsymbol{\Pi}^{\text{new}}=\boldsymbol{\Pi}^{\text{old}}+t^{\text{new}}\left(\overline{\boldsymbol{\Pi}}^{\text{new }}-\boldsymbol{\Pi}^{\text{old}}\right)\vspace*{-0.5em}$ (9)
Improvement with respect to $\boldsymbol{\rho}$ and $\boldsymbol{\Psi}$: For this, we utilize the updated mixing coefficients $\Pi^{\text{new }}$ of the previous step. Since the $\rho_{k}$’s and $\boldsymbol{\psi}_{p,k}$’s are independent for each failure mode, the expression in (8) decouples into $K$ distinct optimization problems. Once the estimators $\rho_{k}^{\text{new }},\psi_{0,k}^{\text{new }}$, and $\boldsymbol{\psi}_{p,k}^{\text{new }}$ are determined, they are used to update the cluster responsibilities in the E-step.
After the EM algorithm is completed, we obtain optimal $\boldsymbol{\Pi}$, $\boldsymbol{\rho}$, and, $\boldsymbol{\Psi}$. We substitute these optimal values to equation (7) to obtain the responsibilities $\boldsymbol{\gamma}_{i}$’s. For each pair of sensor data and failure time, we assign the failure mode with the largest responsibility. This results in class labels for the training set. To perform sensor selection for failure mode $k$, we compute $\left\|\boldsymbol{\psi}_{p,k}\right\|_{2}$, the $l_{2}$-norm of the regression coefficients for sensor $p$ and failure mode $k$. We deem a sensor informative for failure mode $k$ if the $l_{2}$-norm of its regression coefficients is greater than 0.001. The detailed step-wise implementation of the offline sensor selection step is given in algorithm 1.

### v3

@@ 4 Offline Sensor Selection
We consider a scenario where the deep space habitat is monitored by $P$ different sensors. We assume there exists a historical dataset comprised of $P$-dimensional degradation-based sensor signals from monitoring $N$ similar samples. We denote $s_{i,p}(t)$ as the signal data from sensor $p\in\{1,...,P\}$ for system $i\in\{1,...,N\}$ observed at time $t\in[0,T]$, where $T=min(\{TTF_{i}\}_{i=1}^{N})$. By truncating all signals to the minimum failure time, we ensure that all signals in the training set have the same time domain and that they are all included when modeling. For this framework, we assume that the total number of failure modes, $K$ is known, but that the actual failure modes in the training dataset are unknown. This poses a challenge for CA-FPCA, which requires observed covariates for computation. To determine these covariates, we perform a sensor-wise clustering of the training samples. For each sensor, we perform FPCA for feature extraction and then use K-means clustering with $K$ clusters to label the signals from that sensor. Once the sensor-wise clustering is complete, the next step involves feature extraction by applying CA-FPCA and then regressing these features against the time-to-failure to identify the most informative subset of sensors for each failure mode. The next two subsections discuss details about the feature extraction (CA-FPCA), and the optimization (MGR-ASGL) model.
@@ 4.1 Feature Extraction: CA-FPCA
where $v_{i,p}^{k}(t)$ is a smooth random function, and $\epsilon_{i,p}(t)$ are assumed to be independently and identically distributed (i.i.d.) errors with mean zero and variance $\varsigma_{p}^{2}$. $v_{i,p}^{k}(t)$ and $\epsilon_{i,p}(t)$ are assumed to be independent of each other. The mean and covariance functions of $v_{i,p}^{k}(t)$ are given by $\mu_{p}^{k}(t)$ and $\mathbb{C}_{p}^{k}(t,t^{\prime})$, respectively. Using Mercer’s theorem [49], the covariance can be decomposed as follows, $\mathbb{C}_{p}^{k}(t,t^{\prime})=\sum_{m=1}^{\infty}\lambda_{m,p}^{k}\phi_{m,p}^{k}(t)\phi_{m,p}^{k}(t^{\prime})$, where $\{\phi_{m,p}^{k}(t)\}_{m=1}^{\infty}$ are the orthogonal eigenfunctions of $\mathbb{C}_{p}^{k}(t,t^{\prime})$ and $\lambda_{1,p}^{k}\geq...\geq\lambda_{m,p}^{k}\geq...$ are the corresponding eigenvalues. By projecting the mean-subtracted signal data onto the eigenfunctions, we can represent the signals by Eq (2).
@@
To ensure all features are scaled equally, we standardize the CA-FPC-scores by subtracting the sample mean and dividing by the sample standard deviation. Let $x_{i,m}^{p}$ denote the standardized $m$th CA-FPC-score from sensor $p$ of system $i$. We extract $\boldsymbol{x}_{i,p}=(x_{i,1}^{p},...,x_{i,q_{p}}^{p})^{T}$ for sensor $p$ of system $i$,. We then utilize the $\boldsymbol{x}_{i,p}$’s as predictors for fitting the MGR-ASGL model.
@@ 4.2 Optimization of the MGR-ASGL model
Let $Y_{i}=\frac{\ln{TTF}_{i}-(1/N)\sum_{i=1}^{N}\ln{TTF}_{i}}{(1/(N-1))\sum_{i=1}^{N}\big{(}\ln{TTF}_{i}-(1/N)\sum_{i=1}^{N}\ln{TTF}_{i}\big{)}^{2}}$, for $i=1,2…,N$ denote the random variable corresponding to the standardized natural logarithm of the time-to-failure (i.e., $\ln{TTF}$) of system $i$. Furthermore, let $\boldsymbol{Z}_{i}$ denote a categorical random vector of size $K$ used to encode the unknown failure mode of system $i$. If failure mode $k$ is responsible for the failure of system $i$, which occurs with probability $\pi_{k}$, then $\boldsymbol{Z}_{i}[k]=1$ and the remaining elements are zero. In our model, we posit that the relationship between $Y_{i}$ and the observed features $x_{i}$ depends on which failure mode caused the failure of system $i$. Thus, we assume that the conditional distribution of $Y_{i}$ given $\boldsymbol{Z}_{i}[k]=1$ is Gaussian with PDF:
$f_{Y_{i}|\boldsymbol{Z}_{i}}(y_{i}|\boldsymbol{Z}_{i}[k]=1)=\frac{1}{\sqrt{2\pi}\sigma_{k}}\exp\Big{(}-\frac{1}{2\sigma_{k}^{2}}(y_{i}-\beta_{0,k}-\sum_{p=1}^{P}\boldsymbol{x}_{i,p}^{T}\boldsymbol{\beta}_{p,k})^{2}\Big{)}$ (3)
To find the marginal distribution of $Y_{i}$ we use the law of total probability, summing over the joint distribution of $Y_{i}$ and $\boldsymbol{Z}_{i}$ as follows:
$f_{Y_{i}}(y_{i})=\sum_{k=1}^{K}\pi_{k}\frac{1}{\sqrt{2\pi}\sigma_{k}}\exp\Big{(}-\frac{1}{2\sigma_{k}^{2}}(y_{i}-\beta_{0,k}-\sum_{p=1}^{P}\boldsymbol{x}_{i,p}^{T}\boldsymbol{\beta}_{p,k})^{2}\Big{)}$ (4)
Therefore, the marginal distribution of $Y_{i}$ is a mixture of Gaussian regressions. The parameters of this model are $\Theta=\{\pi_{k},\beta_{0,k},\boldsymbol{\beta}_{1,k},\dots,\boldsymbol{\beta}_{P,k},\sigma_{k}\}_{k=1}^{K}$. To estimate these parameters, we seek to minimize the negative incomplete-data log-likelihood (IDLL). To ensure both scale-invariance of parameter estimation and convexity of the optimization problem in the M-step of our EM algorithm, we let $\varphi_{0,k}=\sigma_{k}^{-1}\beta_{0,k}$, $\boldsymbol{\varphi}_{p,k}=\sigma_{k}^{-1}\boldsymbol{\beta}_{p,k}$, and $\rho_{k}=\sigma_{k}^{-1}$ (see: [41]). Thus, the parameters are now $\Theta=\{\boldsymbol{\pi},\boldsymbol{\rho},\boldsymbol{\varphi}\}$, where $\boldsymbol{\pi}=\{\pi_{k}\}_{k=1}^{K}$, $\boldsymbol{\rho}=\{\rho_{k}\}_{k=1}^{K}$, and $\boldsymbol{\varphi}=\{\varphi_{0,k},\boldsymbol{\varphi}_{1,k},\dots,\boldsymbol{\varphi}_{P,k}\}_{k=1}^{K}$. Assuming that $Y_{1},\dots,Y_{N}$ are independent, the negative IDLL is computed as follows:
$-\ell(\Theta|\boldsymbol{Y})=-\sum_{i=1}^{N}\ln\Big{[}\sum_{k=1}^{K}\pi_{k}\rho_{k}\frac{1}{\sqrt{2\pi}}\exp\Big{(}-\frac{1}{2}(y_{i}\rho_{k}-\varphi_{0,k}-\sum_{p=1}^{P}\boldsymbol{x}_{i,p}^{T}\boldsymbol{\varphi}_{p,k})^{2}\Big{)}\Big{]}$ (5)
Estimating $\Theta$ requires minimizing Eq 5. However, the logarithm inside the summation makes Eq 5 difficult to optimize globally. Instead, we search for a local minimum using the EM algorithm, an iterative algorithm suitable for fitting probability models with latent variables. To utilize the EM algorithm, we first construct the negative complete-data log likelihood (CDLL).
@@
$\begin{split}-\ell_{C}(\Theta|\boldsymbol{Y},\mathbb{Z})&=-\ln\prod_{i=1}^{N}\prod_{k=1}^{K}f_{Y_{i}}(y_{i})^{\boldsymbol{Z}_{i}[k]}=-\sum_{i=1}^{N}\sum_{k=1}^{K}\boldsymbol{Z}_{i}[k]\ln\Big{[}\pi_{k}\rho_{k}\frac{1}{\sqrt{2\pi}}\\ &\times\exp\Big{(}-\frac{1}{2}(y_{i}\rho_{k}-\varphi_{0,k}-\sum_{p=1}^{P}\boldsymbol{x}_{i,p}^{T}\boldsymbol{\varphi}_{p,k})^{2}\Big{)}\Big{]}\end{split}$ (6)
In Appendix A, we show that the expectation of the negative CDLL with respect to the conditional distribution of $\mathbb{Z}$ given $\boldsymbol{Y}$ forms an upper bound for the negative IDLL. Therefore, we can obtain a local minimum by minimizing this upper bound. The proposed EM algorithm is performed as follows:
  1. 1.
    Initialization: Set the initial parameters to $\Theta^{(0)}$ s.t.,
    $\Theta^{(0)}=\{\pi_{k}^{(0)},\varphi_{0,k}^{(0)},\boldsymbol{\varphi}_{1,k}^{(0)},\dots,\boldsymbol{\varphi}_{P,k}^{(0)},\rho_{k}^{(0)}\}_{k=1}^{K}$ (7)
  2. 2.
    E-step: Given $\Theta^{(m)}$, the expectation of $-\ell_{C}(\Theta|\boldsymbol{Y},\mathbb{Z})$ is given as:
    $\begin{split}Q(\Theta,\Theta^{(m)})&=\mathbb{E}[-\ell_{C}(\Theta|\boldsymbol{Y},\mathbb{Z})|\boldsymbol{Y},\Theta^{(m)}]=-\sum_{i=1}^{N}\sum_{k=1}^{K}\gamma_{i,k}\ln\Big{[}\pi_{k}\rho_{k}\frac{1}{\sqrt{2\pi}}\\ &\times\exp\Big{(}-\frac{1}{2}(y_{i}\rho_{k}-\varphi_{0,k}-\sum_{p=1}^{P}\boldsymbol{x}_{i,p}^{T}\boldsymbol{\varphi}_{p,k})^{2}\Big{)}\Big{]}\end{split}$ (8)
@@
    $\begin{split}\gamma_{i,k}&=\mathbb{E}(\boldsymbol{Z}_{i}[k]|Y_{i}=y_{i},\Theta^{(m)})=\Pr(\boldsymbol{Z}_{i}[k]=1|Y_{i}=y_{i},\Theta^{(m)})\\ &=\frac{\pi_{k}^{(m)}\rho_{k}^{(m)}\frac{1}{\sqrt{2\pi}}\exp\Big{(}-\frac{1}{2}(y_{i}\rho_{k}^{(m)}-\varphi_{0,k}^{(m)}-\sum_{p=1}^{P}\boldsymbol{x}_{i,p}^{T}\boldsymbol{\varphi}_{p,k}^{(m)})^{2}\Big{)}}{\sum_{l=1}^{K}\pi_{l}^{(m)}\rho_{l}^{(m)}\frac{1}{\sqrt{2\pi}}\exp\Big{(}-\frac{1}{2}(y_{i}\rho_{l}^{(m)}-\varphi_{0,l}^{(m)}-\sum_{p=1}^{P}\boldsymbol{x}_{i,p}^{T}\boldsymbol{\varphi}_{p,l}^{(m)})^{2}\Big{)}}\end{split}$ (9)
    is the responsibility that failure mode $k$ takes for the failure of system $i$. These responsibilities enable a soft clustering, where we claim system $i$ failed due to failure mode $k$ if $\gamma_{i,k}>\gamma_{i,l}$ for $l\neq k$.
  3. 3.
    M-Step: After computing the responsibilities $\gamma_{i,k}$ , we can update our estimate of $\Theta$ by maximizing $Q(\Theta,\Theta^{(m)})$ with respect to $\Theta$. To perform sensor selection, we augment $Q(\Theta,\Theta^{(m)})$ with the adaptive sparse group lasso penalty and form the following optimization problem:
    $\min_{\Theta}\Big{\{}Q(\Theta,\Theta^{(m)})+\lambda\sum_{k=1}^{K}\pi_{k}\Big{(}\alpha\sum_{p=1}^{P}\|\boldsymbol{\varphi}_{p,k}\|_{1}+(1-\alpha)\sum_{p=1}^{P}\sqrt{q_{p}}\|\boldsymbol{\varphi}_{p,k}\|_{2}\Big{)}\Big{\}}$ (10)
    We refer to the problem in Eq 10 as Mixture of Gaussian Regression–Adaptive Sparse Group Lasso (MGR-ASGL). Because $\boldsymbol{\pi}$ is included in the penalty, we follow [51] and use a two-step optimization:
    • (a)
      Step 1: Improvement with respect to $\boldsymbol{\pi}$
      We start by using the responsibilities from the E-step to identify a feasible solution $\bar{\boldsymbol{\pi}}^{(m+1)}=\big{(}\frac{1}{N}\sum_{i=1}^{N}\gamma_{i,1},\dots,\frac{1}{N}\sum_{i=1}^{N}\gamma_{i,K}\big{)}^{T}$. Then, we update the mixing coefficients as follows:
      $\boldsymbol{\pi}^{(m+1)}=\boldsymbol{\pi}^{(m)}+u_{m}(\bar{\boldsymbol{\pi}}^{(m+1)}-\boldsymbol{\pi}^{(m)})$ (11)
      where $u_{m}\in(0,1]$ is chosen to be the largest value in the sequence $\{1,\frac{1}{2},\frac{1}{4},\dots\}$ such that the objective function does not increase.
    • (b)
      Step 2: Improvement with respect to $\boldsymbol{\rho}$ and $\boldsymbol{\varphi}$
      Given the updated mixing coefficients, we can optimize Eq 10 with respect to the other parameters. It can be shown that this problem can be decomposed into solving $K$ individual problems, one for each failure mode. Therefore, for $k=1,\dots,K$, we solve
      $\begin{split}\min_{\rho_{k},\varphi_{0,k},\boldsymbol{\varphi}_{1,k},\dots,\boldsymbol{\varphi}_{P,k}}&\Big{\{}-\sum_{i=1}^{N}\gamma_{i,k}\big{[}\ln\pi_{k}^{(m+1)}+\ln\rho_{k}-\frac{1}{2}(y_{i}\rho_{k}-\varphi_{0,k}\\ &-\sum_{p=1}^{P}\boldsymbol{x}_{i,p}^{T}\boldsymbol{\varphi}_{p,k})^{2}\big{]}\ +\lambda\pi_{k}^{(m+1)}\Big{(}\alpha\sum_{p=1}^{P}\|\boldsymbol{\varphi}_{p,k}\|_{1}\\ &+(1-\alpha)\sum_{p=1}^{P}\sqrt{q_{p}}\|\boldsymbol{\varphi}_{p,k}\|_{2}\Big{)}\Big{\}}\end{split}$ (12)
      Note that the tuning parameters $\lambda$ and $\alpha$ do not change for different $k$. These parameters can be selected using cross-validation. When $\lambda$ increases, ASGL drives groups of regression coefficients to zero. Since these groups correspond to features extracted from individual sensors, this, in effect, causes the influence of some sensors to vanish. Therefore, a sensor is considered significant for failure mode $k$ if $\|\varphi_{p,k}\|_{2}\neq 0$. The parameter $\alpha$ allows for some of the coefficients in a group to drop to zero to improve prediction accuracy. Finally, incorporating the adaptive parameters $q_{p}$ and $\gamma_{i,k}$ into the penalty ensures that sensor selection is not affected by class imbalance and the dimensionality of the features. Let $\boldsymbol{\rho}^{(m+1)}$ and $\boldsymbol{\varphi}^{(m+1)}$ denote the solution to Eq 12. Thus, after completing Steps 1 and 2, we have our updated solution $\Theta^{(m+1)}=\{\boldsymbol{\pi}^{(m+1)},\boldsymbol{\rho}^{(m+1)},\boldsymbol{\varphi}^{(m+1)}\}$.
    Set $\Theta^{(m+1)}\leftarrow\Theta$ and return to (1). Perform successive iterations of (1)(3) until convergence.